# Maximal ideals of reduced group C*-algebras and Thompson’s groups

### 20 Mar 2024 v1

@@ 1. Introduction
We also strengthen [HO17, Theorem 4.5] by showing that, if $F$ is amenable, then $C^{*}_{r}(T)$ has infinitely many ideals (Theorem 4.8).
@@ 3. Reduced group C*-algebras
This is a direct consequence of Theorem 2.11 combined with [BKKO17, Lemma 7.9]. ∎
@@
Let $G$ be a C*-simple group and $H\leq G$ a nontrivial amenable subgroup (for example, $H=\langle g\rangle$ for some $g\in G\setminus\{e\}$). Then, the amenable radical of $K\coloneqq\mathbb{Z}_{2}\wr_{G/H}G$ is $\bigoplus_{G/H}\mathbb{Z}_{2}$, and $\frac{K}{R(K)}=G$ is C*-simple. Thus, a version of Proposition 3.4 (and of Theorem 3.1) does not hold for primitive ideals.
@@ 4. Application: Thompson’s groups
By Lemma 4.1, there exist piecewise linear bijections $\varphi_{1}\colon[y_{1},a]\to[y_{1},\textup{max}\{b,f\}]$ and $\varphi_{2}\colon[a,y_{2}]\to[\textup{max}\{b,f\},y_{2}]$ with finitely many breakpoints in $\mathbb{Z}[1/2]$, and slopes in $2^{\mathbb{Z}}$. Let $g\colon\frac{\mathbb{R}}{\mathbb{Z}}\to\frac{\mathbb{R}}{\mathbb{Z}}$ be given by
@@
The C*-algebra $C^{*}_{\lambda_{T/F}}(T)$ is simple and does not have any tracial states by [KS22, Theorem 6.1]. Therefore, if $F$ is amenable, then the unique nonzero simple quotient of $C^{*}_{r}(T)$ is $C^{*}_{\lambda_{T/F}}(T)$.
@@
Theorems 4.6 and 4.8, together with Remark 4.7, [HO17, Theorem 4.5] and [LBMB18, Corollary 4.2], yield the following characterizations of amenability of $F$ in terms of the ideal structure of $C^{*}_{r}(T)$.

### 22 Mar 2024 v2

@@ 1. Introduction
We also strengthen the result of Haagerup and Olesen ([HO17, Theorem 5.5]) by showing that, if $F$ is amenable, then $C^{*}_{r}(T)$ has infinitely many ideals (Theorem 4.8).
@@ 3. Reduced group C*-algebras
This is a direct consequence of Theorem 2.11 and [BKKO17, Lemma 7.9]. ∎
@@
Let $G$ be a C*-simple group and let $H\leq G$ be a nontrivial amenable subgroup (for example, $H=\langle g\rangle$ for some $g\in G\setminus\{e\}$). Then, the amenable radical of $K\coloneqq\mathbb{Z}_{2}\wr_{G/H}G$ is $\bigoplus_{G/H}\mathbb{Z}_{2}$, and $\frac{K}{R(K)}=G$ is C*-simple. Thus, a version of Proposition 3.4 (and of Theorem 3.1) does not hold for primitive ideals.
@@ 4. Application: Thompson’s groups
By Lemma 4.1, there exist piecewise linear bijections
$\varphi_{1}\colon[y_{1},a]\to[y_{1},\textup{max}\{b,f\}]$
and
$\varphi_{2}\colon[a,y_{2}]\to[\textup{max}\{b,f\},y_{2}]$
with finitely many breakpoints in $\mathbb{Z}[1/2]$, and slopes in $2^{\mathbb{Z}}$. Let $g\colon\frac{\mathbb{R}}{\mathbb{Z}}\to\frac{\mathbb{R}}{\mathbb{Z}}$ be given by
@@
The C*-algebra $C^{*}_{\lambda_{T/F}}(T)$ is simple and does not have any tracial states by [KS22, Theorem 6.1]. It follows from Theorem 4.6 and the fact that $C^{*}_{\lambda_{T/F}}(T)$ is simple that, if $F$ is amenable, then the unique nonzero simple quotient of $C^{*}_{r}(T)$ is $C^{*}_{\lambda_{T/F}}(T)$.
@@
Theorems 4.6 and 4.8, together with Remark 4.7, [HO17, Theorem 5.5] and [LBMB18, Corollary 4.2], yield the following characterizations of the amenability of $F$ in terms of the ideal structure of $C^{*}_{r}(T)$.